Suppose that a shuttles launch depends on three key devices that operate independently
of each other and malfunction with probabilities 0.01, 0.02, and 0.02, respectively. If any of the key devices
malfunctions, the launch will be postponed. Compute the probability for the shuttle to be launched on time,
according to its schedule.

Answer by ikleyn(52847)

.

P = (1-0.01)*(1-0.02)*(1-0.02) = 0.99*0.98*0.98.    


You compute the value.